# PowerPlots: An Open Source Power Grid Visualization and Data Analysis Framework for Academic Research

### v1

@@ Abstract
Data visualization is important for developing an understanding of a complex system. PowerPlots.jl is a data visualization tool for power grids, one of the most complex systems in the world. The design of PowerPlots.jl is intended to facilitate exploration of power grid data while performing research and to facilitate communication of research findings to an audience. Several tools created to support this software also facilitate analysis of power grid data by transforming the data into graph topology or data-frame data formats that are more compatible for some applications. The high level of flexibility in PowerPlots.jl enables researchers who are developing and analyzing methods for solving novel power grid problems to better understand and communicate the complexities of their research.
@@ I Introduction
The power grid is a complex engineered network system and its operation is governed by the physics of power flow. Developing an intuition for the operation of the power grid is a valuable skill that allows a researcher to identify when a unique behavior is occurring or more easily identify an error in their computational method. Power grid visualization allows the researcher to see and recognize these patterns [1].
Visualization of the power grid has two primary values. 1) Exploration: It allows for exploration of data and debugging of methods by quickly and interactively visualizing data, 2) Communication: it enables sharing of specific information to communicate research findings. The two goals of PowerPlots.jl are to enable the exploration of a research problem and the communication of research results.
@@
Many software packages, libraries, or ecosystems exist that can also perform power grid visualization, but none cohesively possess all of the additional flexible capabilities needed for supporting research applications, such as the ability to view and explore custom data, extend data formats with new information, and extensively customize figures for publication. PyPSA [2] or Python for Power System Analysis is an open source energy system modeling project. It supports a range of standard power grid optimization models like unit commitment and capacity expansion, and includes cross sector energy systems modeling. It includes plotting of networks for its users to visualize the results and view the solutions of these models. OpenDSS [3] is a distribution network simulator developed by EPRI. Originally a scripting tool to define networks and routines for simulation, an extension for a graphical environment was developed, OpenDSS-G, to enable a visual interface to create and view simulation results. PowerGridModel [4] is a an open source power grid analysis library that supports power flow, state estimation, and short circuit analysis. It includes visualization software in its data analysis extension [5]. PandaPower [6] is an open source python package for power system analysis. It emphasizes a tabular data structure for data analysis of power flow, state estimation, and short-circuit calculations. It has basic support for visualizing the network representation of the data.
@@
PowerWorld [7] is an industry software initially developed to provide students with a visual tool to understand the power system concepts that they learned in classes. The slogan “the visual approach to electric power systems” emphasizes its roots as a tool to develop intuition outside of looking at charts and tables of power grid data. A variety of other commercial industry software including PSSE [8], PLSF [9], ETAP [10], and TARA [11] have some level of visualization of network data like prices or outage reliability. These tools typically have very limited flexibility to visualize data beyond what they are designed to calculate.
@@
PowerPlots.jl is a flexible tool for creating visualizations of power grids to better support the needs of researchers. It has a simple interface for creating a network plot to view and interact with data, and a high level of customization to create publication-ready figures. It is written in the Julia programming language [12], and adopts the data format from PowerModels.jl [13] and PowerModelsDistribution.jl [14], which parse file formats including MatPower [15], PSSE [8], as well as a JSON dictionary. The data format is extendable to visualize new components that are not natively supported by these file formats. In addition, several tools like a graph data model and a data-frame data model, created to support the network plots, are also useful as data analysis tools for research.
Many grid research packages can take advantage of PowerPlots.jl to enable analysis of their outputs with no change to their code [16, 17, 18, 19], and many publications already use PowerPlots.jl to communicate their results in application areas ranging from wildfire risk, to carbon intensity metrics, restoration planning, and network reconfiguration  [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
This paper is not documentation for PowerPlots.jl; the documentation can accessed at the GitHub repository of the package [^1] https://wispo-pop.github.io/PowerPlots.jl/dev/. Rather, this paper explores how the package can be used as a research tool, including examples of the types of analyses that this package enables.
The primary contributions of this paper are (1) An introduction to the plotting tool PowerPlots.jl and its software design, (2) Examples of how to use PowerPlots.jl to explore network data and customize figures, and (3) An exploration of the types of data analysis that are provided by tools within PowerPlots.jl.
The rest of the paper is organized as follows: Section II explains the design and basic software structure of PowerPlots.jl. Section III describes the features of PowerPlots.jl to enable data exploration, with examples of how to investigate different types of research questions. Section III-F demonstrates the customization that can be achieved using PowerPlots.jl. Section IV shows how the data structures within PowerPlots.jl can be utilized to conduct data analyses of power grid data. Finally, Section V concludes the work.
@@ V Conclusion
Visualization of data is a powerful tool for building an understanding of data and highlighting aspects for communication. PowerPots.jl has a simple interface to construct a view of the network information, but has significant flexibility to enable visualization of any parameter of interest. It also allows significant customization to enable a user to create a figure that highlights the important information they wish to communicate. In creating PowerPlots.jl, several tools are developed that also provide utility in data analysis of the power grid. These include data transformations to graph network models and to tabular data formats that make applicable data analyses easier to perform. Together, the visualization and analysis tools in PowerPlots.jl enable researchers to explore novel power grid problems by providing capabilities that support research beyond standard power grid data. This software package will continue to evolve and develop new features. Additional customization features will be available from the user arguments, and the implementation of these features will be guided by community involvement in this package. In addition, future extensions of this work can be applied to visualizing other network systems such as gas, water, or telecommunications networks. This will eventually enable a visual understanding for cross-sector critical infrastructure analysis.
@@ Acknowledgments
I would like to thank my PhD advisor Line Roald for her encouragement in developing this package, Hanbin Yang, Kshitij Girigoudar, and Joe Gorka for being early users of this package, and the people who have participated in the open source project by reporting issues and giving suggestions online. I would also like to thank the Advanced Network Science Initiative at Los Alamos National Laboratory for creating the InfrastructureModels.jl ecosystem on which PowerPlots.jl was developed. Support for this work was provided in part by the National Science Foundation (NSF) under the NSF CAREER award No. 2045860 and NSF ASCENT award No. 2132904, and by the Department of Energy through the LANL/LDRD Program and the Center for Nonlinear Studies.

### v2

# PowerPlots.jl: An Open Source Power Grid Visualization and Data Analysis Framework for Academic Research

@@ Abstract
Data visualization is essential for developing an understanding of a complex system. The power grid is one of the most complex systems in the world and effective power grid research visualization software must 1) be easy to use, 2) support unique data that may arise in research, and 3) be capable of creating custom figures for publication and presentation. However, no current software addresses all three of these needs. PowerPlots is an open-source data visualization tool for power grids that does address these needs. In addition, several tools created to support this software facilitate the analysis of power grid data by transforming the data into graph topology or data-frame data formats that are more compatible for some analyses. In this work, we use PowerPlots to investigate several case studies that involve exploring power grid data. These case studies demonstrate the valuable insights that are possible when using network visualization and how it can be applied to research applications.
Index Terms:
Visualization, Power Grid, Network, Data Analysis, Julia Langauge
@@ I Introduction
The power grid is a complex engineered network system and its operation is governed by the physics of power flow. Developing an intuition for the operation of the power grid is a valuable skill that allows a researcher to identify when a unique behavior is occurring or more easily identify an error in their computational method. Power grid visualization allows the researcher to see and recognize these patterns [1]. Visualization of the power grid gives intuition through two use-cases. 1) Exploration: it allows for exploration of data and debugging of methods by quickly and interactively visualizing data, and 2) Communication: it enables sharing of specific information to communicate research findings.
To support these use-cases, visualization software must be easy to use, support unique data that may arise in research, and be capable of creating custom figures for publication and presentation. Many packages, libraries, or ecosystems exist that can perform power grid visualization, but none have all three of these traits. As a result, they may not be able to view and explore custom data, extend data formats with new information, and extensively customize figures for publication.
@@
Several existing open-source power grid modeling packages have a limited support for network visualization. They can show basic network structure and possibly show traditional power study data, but not unique research data from a novel power grid problem. PyPSA [2] or Python for Power System Analysis is an open source energy system modeling project. It supports a range of standard power grid optimization models like unit commitment and capacity expansion, and includes cross sector energy systems modeling. It includes plotting of networks for its users to visualize the results and view the solutions of these models. OpenDSS [3] is a distribution network simulator developed by EPRI. Originally a scripting tool to define networks and routines for simulation, an extension for a graphical environment was developed, OpenDSS-G, to enable a visual interface to create and view simulation results. PowerGridModel [4] is a an open source power grid analysis library that supports power flow, state estimation, and short circuit analysis. It includes visualization software in its data analysis extension [5]. PandaPower [6] is an open source python package for power system analysis. It emphasizes a tabular data structure for data analysis of power flow, state estimation, and short-circuit calculations. It has basic support for visualizing the network representation of the data.
@@
Many closed-source software ecosystems are even more limited in their ability to show unique research data. These tools typically have very limited flexibility to visualize data beyond what they are designed to calculate. PowerWorld [7] is an industry software initially developed to provide students with a visual tool to understand the power system concepts that they learned in classes. The slogan “the visual approach to electric power systems” emphasizes its roots as a tool to develop intuition outside of looking at charts and tables of power grid data. A variety of other commercial industry software including PSSE [8], PLSF [9], ETAP [10], and TARA [11] have some level of visualization of network data like prices or outage reliability.
@@
PowerPlots.jl is a flexible tool for creating visualizations of power grids to better support the needs of researchers. It has a simple interface for creating a network plot to view and interact with data, supports custom data fields, and has a high level of customization to create publication-ready figures. It is written in the Julia programming language [12], and adopts the data format from PowerModels.jl [13] and PowerModelsDistribution.jl [14], which parse file formats including MatPower [15], PSSE [8], as well as a JSON dictionary. The data format is extendable to visualize new components that are not natively supported by these file formats. Furthermore, several tools like a graph data model and a data-frame data model, created to support the network plots, are also useful as data analysis tools for research.
Many grid research packages can take advantage of PowerPlots.jl to enable analysis of their outputs with no change to their code [16, 17, 18, 19], and many publications already use PowerPlots.jl to communicate their results in application areas including wildfire risk, carbon intensity metrics, restoration planning, and network reconfiguration  [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
This paper is not documentation for PowerPlots.jl; the documentation can be accessed at the GitHub repository of the package [^1] https://wispo-pop.github.io/PowerPlots.jl/dev/. Rather, this paper explores how the package can be used as a research tool, including examples of the types of analyses that this package enables.
The primary contributions of this paper are (1) An introduction to the plotting tool PowerPlots.jl and its software design, (2) Several case studies that use PowerPlots.jl to explore network data and customize figures, and (3) An exploration of the types of data analysis that are provided by tools within PowerPlots.jl.
The rest of the paper is organized as follows: Section II explains the design and basic software structure of PowerPlots.jl. Section III describes the features of PowerPlots.jl to enable data exploration, with case studies of how to investigate different types of research questions. Section IV shows how the data structures within PowerPlots.jl can be utilized to conduct data analyses of power grid data. Finally, Section V concludes the work.
@@ V Conclusion
Visualization of data is a powerful tool for building an understanding of data and highlighting aspects for communication. PowerPots.jl has a simple interface to construct a view of the network information, but has significant flexibility to enable visualization of any parameter of interest. It also allows significant customization to enable a user to create a figure that highlights the important information they wish to communicate. In creating PowerPlots.jl, several tools are developed that also provide utility in data analysis of the power grid. These include data transformations to graph network models and to tabular data formats that make applicable data analyses easier to perform. Together, the visualization and analysis tools in PowerPlots.jl enable researchers to explore novel power grid problems by providing capabilities that support research beyond standard power grid data. This software package will continue to evolve and develop new features. Additional customization features will be available from the user arguments, and the implementation of these features will be guided by community involvement in this package. Furthermore, future extensions of this work can be applied to visualizing other network systems such as gas, water, or telecommunications networks. This will eventually enable a visual understanding for cross-sector critical infrastructure analysis.
@@ Acknowledgments
I would like to thank my PhD advisor Line Roald for her encouragement in developing this package, Hanbin Yang, Kshitij Girigoudar, and Joe Gorka for being early users of this package, and the people who have participated in the open source project by reporting issues and giving suggestions online. I would also like to thank the Advanced Network Science Initiative at Los Alamos National Laboratory for creating the InfrastructureModels.jl ecosystem on which PowerPlots.jl was developed.